OpenStudy (anonymous):

Seven inches are equivalent to 17.78 cm. Find the equation converting length in cm to length in inches. What is the length in inches of a 96.2 cm bar

OpenStudy (anonymous):

1 inch = 2.54 cm x inches = 96.2 x = 37.84 inches

Directrix (directrix):

Inches is to 7 inches as c centimeters is to 17.78 cm I/7 = c/17.78 17.78 I = 7 c I = 7c/17.78 I(c) = 7/17.78 (c) where c is the length given in centimeters and I is the conversion measure in inches --> Equation I(96.2) = (7/17.78) * (96.2) I = 37.874 inches (approx)